# On Stealing Graph Neural Network Models

### v1

@@ Abstract
Current graph neural network (GNN) model-stealing methods rely heavily on queries to the victim model, assuming no hard query limits. However, in reality, the number of allowed queries can be severely limited. In this paper, we demonstrate how an adversary can extract the GNN with very limited interactions with the model. Our approach first enables the adversary to obtain the model backbone without making direct queries to the victim model and then to strategically utilize a fixed query limit to extract the most informative data. The experiments on eight real-world datasets demonstrate the effectiveness of the attack, even under a very restricted query limit and under defense against model extraction in place. Our findings underscore the need for robust defenses against GNN model extraction threats.
@@ 1 Introduction
Collectively, our results demonstrate that model-stealing attacks on GNNs are both more effective and resource-efficient than previously understood, highlighting the critical severity of these attacks in both inductive and transductive scenarios. For instance, targeting a SAGE (hamilton2017inductive) model trained on the Physics dataset, we achieve 91% accuracy with only 100 queries to the victim model –— compared to approximately 5,000 queries, higher computational cost and additional victim output (such as embeddings) needed by the current state-of-the-art method to reach similar accuracy.
@@ 2.1 Graph Neural Networks
Our work addresses model-stealing vulnerabilities in GNNs, specifically for node-level tasks. For such tasks, GNNs are typically trained and evaluated in two primary scenarios crucial to understanding these vulnerabilities: transductive and inductive settings. In the transductive paradigm, the data consists of a fixed graph with a portion of nodes labeled for training, while another portion remains unlabeled. The objective is to apply these labeled nodes to predict labels for the previously unlabeled ones within the same graph. However, transductive models are limited in their ability to generalize to novel nodes. Conversely, in the inductive setting, GNNs are designed to expand their learning to accommodate new, unseen nodes or graphs that were not part of the training data.
@@ 2.2 Model Stealing Attacks
| Trans. | Ind. | Data Access | Victim Q. Access | Victim Output | |
@@
| Podhajski_2024 | ✗ | ✓ | Limited | Unlimited | Embed. and Pred. |
| shen2021model | ✗ | ✓ | Limited | Unlimited | Embed. and Pred. |
@@
Relation to existing work. An adversary attempting to steal a model faces two primary challenges: limited data access or restricted access to the model itself. Existing approaches to GNN model stealing predominantly address scenarios with limited or no data access, but none explore the problem of severely restricted model access. Specifically, shen2021model and Podhajski_2024 concentrate on stealing the encoder in the inductive setting, assuming unlimited access to the model but limited access to data. Both of these methods leverage the victim model’s output (e.g., embeddings) to train the encoder and then train Multi-Layer Perceptrons (MLPs) on top of the frozen encoder using available class labels. This requirement for intermediate embeddings contrasts with our work, which assumes a more restrictive and practical threat model where adversaries only receive the final class predictions. Similarly, wu2021model focuses on limited data access in the transductive setting, while also assuming unrestricted access to the victim model. Another contribution in this domain is the data-free model extraction attack framework proposed by datafree. This framework enables GNN model extraction without requiring access to actual node features or graph structures. However, it assumes unlimited access to the victim model and relies on an extensive number of queries.
@@
2.3 Self-supervised learning in GNN
@@ 2.4 Motivation
Existing works on inductive GNN model stealing, such as those by shen2021model and Podhajski_2024, rely on stealing the victim’s encoder using rich responses like query embeddings. However, a key observation from self-supervised learning provides a more efficient, query-free path for the adversary. In the inductive graph learning setting, velickovic2018deep showed that a randomly initialized graph convolutional network (GCN) can already extract informative features and serve as a strong baseline. This phenomenon is linked to the Weisfeiler-Lehman graph isomorphism test (weisfeiler1968reduction), as discussed by (KW17; hamilton2017inductive). As shown by recent works, a randomly initialized GNN encoder, when paired with a trained MLP head, can achieve performance comparable to that of a fully trained model. For instance, DGI (velickovic2018deep) reports test accuracies of 93.3% (random encoder) vs. 94.0% (SSL-trained) on Reddit, and 62.6% vs. 63.8% on PPI. Similarly, BGRL (thakoor2022large) observes minimal performance gaps across datasets such as WikiCS, Computer, Photo, CS, and Physics. These results are summarized in Table 2. Our experimental findings align with these insights. T-SNE visualizations (Figure 4 in the Appendix) suggest that even randomly initialized encoders produce structured embeddings in the inductive setting.
In contrast, the transductive setting sees greater benefits from self-supervised learning. As shown in Table 2, training a GNN encoder with SSL yields notable performance gains. Since the full graph structure, including test node connectivity, is available during training, the encoder can learn representations that better align with the downstream task—unlike in the inductive setting, where such structural information is unavailable (velickovic2018deep; thakoor2022large). These improvements are further supported by the visualizations in Figure 4 in the Appendix. Notably, transductive graphs are typically small, keeping SSL training costs low and making it a viable option even with limited computational resources.
@@ 3.3 Query Selection
In the final step, we aim to select a data subset that maximizes information extraction from the victim model within the query limit. Intuitively, our goal is to select the most diverse subset possible, ensuring comprehensive coverage of the dataset accessible to the adversary. This strategy of selecting representative points to cover the input space is conceptually analogous to diversity sampling methods in the field of active learning (settles_active_2009). This selected subset is then used to train an MLP head on top of the encoder, enabling the construction of a high-quality surrogate model.
@@ 4.1 Stealing with a Random Encoder and Self-Supervised Learning
The results of this comparison, as presented in Tables 3 and 6 (Appendix), include the accuracy and fidelity of both approaches, reported along with their standard deviations. These evaluations were conducted using randomly selected queries, without leveraging any query selection algorithm. The results reveal that the performance of the frozen encoder with an MLP is comparable to that of the E2E approach across most datasets, as well as when using both GAT and SAGE target models. Additionally, Figures 3 and 5 present charts that confirm these results hold across different query limits. Importantly, the R-init method requires significantly fewer computational resources, as only the MLP layer needs training, whereas the E2E approach involves updating the entire network. We used T-SNE to visualize embeddings from the Physics dataset, as shown in Figure 4. The 2D plot compares embeddings from a random encoder and a trained encoder. Clusters corresponding to node classes are clearly separated, with the random encoder producing well-defined clusters closely resembling those of the trained encoder.
In addition to the inductive setting, we also explore the transductive setting by evaluating the impact of self-supervised learning on encoder performance. Beyond the E2E approach, we consider an SSL-trained encoder paired with an MLP (SSL). The results, summarized in Tables 4 and 7 (in the Appendix), demonstrate a significant improvement in accuracy and fidelity when using SSL compared to relying solely on the query data set. This improvement is particularly notable in datasets such as Cora and Citeseer and is observed consistently across different target models (GCN and GAT) as well as surrogate architectures (GIN, GCN, and SAGE). Figure 6 confirms this result across different query limits, showing that the performance improvement from SSL becomes more pronounced as the query limit decreases. In addition, the findings suggest that incorporating all available adversarial data in local self-supervised training yields substantial performance gains across all benchmarks. This underscores the utility of SSL in extracting high-quality representations, even in extreme scenarios, enhancing the overall robustness and effectiveness of the model-stealing process.
@@ 4.2 Strategic Query Selection
The results show that selecting queries with K-means based on the encoder embeddings results in higher accuracy and fidelity in both setups across all of the datasets and target models. The results also indicate that in the transductive setting, our method provides a selection of more diverse queries based on the SSL-trained embeddings. For the inductive setting, it is further confirmed that the representations produced by the random encoder are meaningful and possible to interpret. By comparing the accuracy of our surrogate models with the performance of the victim models, we observe that the surrogate models achieve comparable results with significantly fewer labeled data. Additionally, Figure 3 and Figures 5 and 6 in the Appendix illustrate how performance improves with different query limits, further demonstrating that as the query limit decreases, the effectiveness of our method increases.
@@
We also conduct McNemar’s test, as presented in Tables 8 and 9 in the Appendix, to compare the stolen model with the original model by evaluating their classification errors on the same dataset. The null hypothesis assumes no significant difference in the classification error rates between the stolen and original models. The results demonstrate that our method produces a stolen model with a higher degree of similarity to the victim model.
| Method | Photo | Physics | ||||
| Acc. | Fid. | Rel. Time | Acc. | Fid. | Rel. Time | |
| Target accuracy | 93.0 | 96.0 | ||||
@@ 4.3 Comparison with existing methods
Inductive setting. We compare our method against all existing approaches, i.e., (shen2021model) and (Podhajski_2024). We note that these two previous approaches do not take into account scenarios where the adversary’s dataset exceeds the query limit. Thus, to explore this scenario, we replicate the experiments described in (shen2021model) and (Podhajski_2024) with randomly selected queries from the entire available dataset. Additionally, we compare our performance with (datafree), which demonstrates that an adversary can successfully steal a model without access to any training data, though this requires a substantial number of queries: 100 queries for graphs of size 250, totaling 25,000 query nodes. We show that our method performs much better in a scenario with very restricted model access (we evaluate our approach on 10, 25, 50, 100, and 500 nodes). We present our results in Table 3 and Appendix Table 6, where we compare all methods using the same surrogate model architecture and different target architectures: SAGE and GAT, with a query limit of 100. We also compare the performance of all methods across different query limits in Figure 3 and Appendix Figure 5. Furthermore, in Table 5, we compare the performance using the architectures originally employed in (shen2021model) and (Podhajski_2024). Our method, which does not rely on query embeddings from the victim model, consistently outperforms the previous methods in both accuracy and fidelity. This superiority is evident across various baseline surrogate models, including GIN, GAT, and SAGE. Additionally, by utilizing the randomly initialized encoder, we significantly improved the execution times as seen in Table 5. Moreover, even when the same architecture is used, our approach yields significantly better results, demonstrating its effectiveness.
@@
Transductive setting. For the transductive setting, there is no previous work in the literature to which we can directly compare. However, we believe that in the work by (wu2021model) we can find a setting to which a comparison of our E2E is reasonable. In particular, unlike our work, (wu2021model) focuses on the scenarios where the adversary lacks some types of data, e.g., the graph structure or node attributes. In contrast, our focus is on scenarios where the adversary is not limited in terms of data type but in terms of the number of queries that can be made. To create a relevant comparison, we modified the approach by (wu2021model) by training their model using randomly selected queries under the assumption that it has access to the entire data available to the adversary (with no restrictions on data types). Similarly, we compare our results to (datafree), which generates graph queries in a restricted scenario, where the query limit is reduced to just 5, 10, 20, or 50 nodes. The results presented in Tables 4 and 7 (in the Appendix) show a significant improvement in both accuracy and fidelity when our method is applied. Specifically, when the surrogate encoder is trained with all available data and the query selection is optimized using the K-means clustering algorithm, the performance of the stolen model is greatly enhanced. This demonstrates the advantage of our approach in transductive settings, where strategic data utilization and query selection can significantly boost the attack’s effectiveness. We further validate the effectiveness of our method across different surrogate architectures (GIN, SAGE, and GCN) and observe that our method consistently outperforms all baselines regardless of the surrogate used. Notably, the GCN surrogate yields the highest performance overall, surpassing all other methods and architectural configurations. Additionally, Figure 6 illustrates how the performance improves with different query limits, further emphasizing the impact of our method as the query limit decreases.
@@ 5 Defense
We evaluate all stealing methods under a defense that flips class predictions with a probability of in both inductive and transductive settings (Tables 10 and 11 in the Appendix). Our proposed method consistently achieves the highest accuracy and fidelity, even in the presence of defense measures. These results highlight the difficulty of defending against our model extraction attack, even under strong defense mechanisms that come at the cost of reducing the utility of the model.
6 Limitations
In this paper, we use the K-means algorithm to select a subset of nodes that covers the embedding space, dividing it into clusters. Although our approach demonstrates an improved performance of the stealing model, further research could explore alternative methods for selecting the query set. Additionally, we show that, in the inductive setting, a randomly initialized encoder achieves performance similar to that of a trained encoder, allowing the adversary to steal the model with fewer resources. Although random encoders perform less competitively in the transductive setting compared to SSL, investigating the potential to enhance model stealing in this context with reduced resources would be a valuable direction for future research.

### v2

@@ Abstract
Current graph neural network (GNN) model-stealing methods rely heavily on queries to the victim model, assuming no hard query limits. However, in reality, the number of allowed queries can be severely limited. In this paper, we demonstrate how an adversary can extract a GNN with very limited interactions with the model. Our approach first enables the adversary to obtain the model backbone without making direct queries to the victim model and then to strategically utilize a fixed query limit to extract the most informative data. The experiments on eight real-world datasets demonstrate the effectiveness of the attack, even under a very restricted query limit and under defense against model extraction in place. Our findings underscore the need for robust defenses against GNN model extraction threats.
@@ 1 Introduction
Collectively, our results demonstrate that model-stealing attacks on GNNs are both more effective and resource-efficient than previously understood, highlighting the critical severity of these attacks in both inductive and transductive scenarios. For instance, targeting a SAGE (hamilton2017inductive) model trained on the Physics dataset, we achieve 91% accuracy with only 100 queries to the victim model –— compared to approximately 5,000 queries, higher computational cost, and additional victim output (such as embeddings) needed by the current state-of-the-art method to reach similar accuracy.
@@ 2.1 Graph Neural Networks
Our work addresses model-stealing vulnerabilities in GNNs, specifically for node-level tasks. For such tasks, GNNs are typically trained and evaluated in two scenarios crucial to understanding these vulnerabilities: transductive and inductive settings. In the transductive paradigm, the data consists of a fixed graph with a portion of nodes labeled for training, while another portion remains unlabeled. The objective is to apply these labeled nodes to predict labels for the previously unlabeled ones within the same graph. However, transductive models are limited in their ability to generalize to novel nodes. Conversely, in the inductive setting, GNNs are designed to expand their learning to accommodate unseen nodes or graphs that were not part of the training data.
@@ 2.2 Model Stealing Attacks
| Trans. | Ind. | Data Acc. | Vict. Q. Acc. | Vict. Output | |
@@
| Podhajski_2024 | ✗ | ✓ | Limited | Unlimited | Emb. & Pred. |
| shen2021model | ✗ | ✓ | Limited | Unlimited | Emb. & Pred. |
@@
Relation to existing work. An adversary attempting to steal a model faces two primary challenges: limited data access or restricted access to the model itself. Existing approaches to GNN model stealing predominantly address scenarios with limited or no data access, but none explore the problem of severely restricted model access. Specifically, shen2021model and Podhajski_2024 concentrate on stealing the encoder in the inductive setting, assuming unlimited access to the model but limited access to data. Both of these methods leverage the victim model’s output (e.g., embeddings) to train the encoder and then train Multi-Layer Perceptrons (MLPs) on top of the frozen encoder using available class labels. This requirement for intermediate embeddings contrasts with our work, which assumes a more restrictive and practical threat model where adversaries only receive the final class predictions. Similarly, wu2021model focuses on limited data access in the transductive setting, while also assuming unrestricted access to the victim. Another contribution is the data-free model extraction attack framework proposed by datafree. This framework enables GNN model extraction without requiring access to actual node features or graph structures. However, it assumes unlimited access to the victim model and relies on an extensive number of queries.
@@
2.3 Self-Supervised Learning in GNNs
@@ 2.4 Motivation
Existing works on inductive GNN model stealing, such as those by shen2021model and Podhajski_2024, rely on stealing the victim’s encoder using rich responses like query embeddings. However, a key observation from self-supervised learning provides a more efficient, query-free path for the adversary. In the inductive graph learning setting, velickovic2018deep showed that a randomly initialized graph convolutional network (GCN) can already extract informative features and serve as a strong baseline. This phenomenon is linked to the Weisfeiler-Lehman graph isomorphism test (weisfeiler1968reduction), as discussed by KW17; hamilton2017inductive. As shown by recent works, a randomly initialized GNN encoder, when paired with a trained MLP head, can achieve performance comparable to that of a fully trained model. For instance, DGI (velickovic2018deep) reports test accuracies of 93.3% (random encoder) vs. 94.0% (SSL-trained) on Reddit, and 62.6% vs. 63.8% on PPI. Similarly, BGRL (thakoor2022large) observes minimal performance gaps across datasets such as WikiCS, Computer, Photo, CS, and Physics. These results are summarized in Table 2. Our experimental findings align with these insights. T-SNE visualizations (Figure 3 in the Appendix) suggest that even randomly initialized encoders produce structured embeddings in the inductive setting.
In contrast, the transductive setting sees greater benefits from self-supervised learning. As shown in Table 2, training a GNN encoder with SSL yields notable performance gains. Since the full graph structure (including test node connectivity) is available during training, the encoder can learn representations that better align with the downstream task – unlike in the inductive setting, where such structural information is unavailable (velickovic2018deep; thakoor2022large). These improvements are further supported by the visualizations in Figure 3 in the Appendix. Notably, transductive graphs are typically small, keeping SSL training costs low and making it a viable option even with limited computational resources.
@@ 3.3 Query Selection
In the final step, we aim to select a data subset that maximizes information extraction from the victim model within the query limit. Intuitively, our goal is to select the most diverse subset possible, ensuring comprehensive coverage of the dataset accessible to the adversary. This strategy of selecting representative points to cover the input space is conceptually analogous to diversity sampling methods in the field of active learning (settles_active_2009; 53258). This selected subset is then used to train an MLP head on top of the encoder, enabling the construction of a high-quality surrogate model.
@@ 4.1 Stealing with a Random Encoder and Self-Supervised Learning
The results of this comparison, as presented in Tables 3 and 5 (Appendix), include the accuracy and fidelity of both approaches, reported along with their standard deviations. These evaluations were conducted using randomly selected queries, without leveraging any query selection algorithm. The results reveal that the performance of the frozen encoder with an MLP is comparable to that of the E2E approach across most datasets, as well as when using both GAT and SAGE target models. Additionally, Figures 4 and 5 (Appendix) present charts that confirm these results hold across different query limits. Importantly, the R-init method requires significantly fewer computational resources, as only the MLP layer needs training, whereas the E2E approach involves updating the entire network. We used T-SNE to visualize embeddings from the Physics dataset, as shown in Figure 3 (Appendix). The 2D plot compares embeddings from a random encoder and a trained encoder. Clusters corresponding to node classes are clearly separated, with the random encoder producing well-defined clusters closely resembling those of the trained encoder.
In addition to the inductive setting, we also explore the transductive setting by evaluating the impact of self-supervised learning on encoder performance. Beyond the E2E approach, we consider an SSL-trained encoder paired with an MLP (SSL). The results, summarized in Tables 4 and 6 (in the Appendix), demonstrate a significant improvement in accuracy and fidelity when using SSL compared to relying solely on the query data set. This improvement is particularly notable in datasets such as Cora and Citeseer and is observed consistently across different target models (GCN and GAT) as well as surrogate architectures (GIN, GCN, and SAGE). Figure 6 (in the Appendix) confirms this result across different query limits, showing that the performance improvement from SSL becomes more pronounced as the query limit decreases. In addition, the findings suggest that incorporating all available adversarial data in local self-supervised training yields substantial performance gains across all benchmarks. This underscores the utility of SSL in extracting high-quality representations, even in extreme scenarios, enhancing the overall robustness and effectiveness of the model-stealing process.
@@ 4.2 Strategic Query Selection
The results show that selecting queries with K-means based on the encoder embeddings results in higher accuracy and fidelity in both setups across all of the datasets and target models. The results also indicate that in the transductive setting, our method provides a selection of more diverse queries based on the SSL-trained embeddings. For the inductive setting, it is further confirmed that the representations produced by the random encoder are meaningful and possible to interpret. By comparing the accuracy of our surrogate models with the performance of the victim models, we observe that the surrogate models achieve comparable results with significantly fewer labeled data. Additionally, Figures 4, 5 and 6 in the Appendix illustrate how performance improves with different query limits, further demonstrating that as the query limit decreases, the effectiveness of our method increases. To justify the use of our approach in the Select phase, we compare K-means with several representative selection strategies, including farthest-first, K-center greedy, entropy sampling, coreset herding, and margin sampling (settles_active_2009; scheffer2001active; welling2009herding; gonzalez1985clustering). As shown in Table 10 in the Appendix, while methods such as coreset herding offer improvements over random selection, K-means consistently delivers the highest accuracy and fidelity across all datasets.
@@
We also conduct McNemar’s test (McNemar1947), as presented in Tables 7 and 8 in the Appendix, to compare the stolen model with the original model by evaluating their classification errors on the same dataset. The null hypothesis assumes no significant difference in the classification error rates between the stolen and original models. The results demonstrate that our method produces a stolen model with a higher degree of similarity to the victim model.
@@ 4.3 Comparison with Existing Methods
Inductive setting. We compare our method against all existing approaches, i.e., shen2021model and Podhajski_2024. We note that these two previous approaches do not take into account scenarios where the adversary’s dataset exceeds the query limit. Thus, to explore this scenario, we replicate the experiments described in shen2021model and Podhajski_2024 with randomly selected queries from the entire available dataset. Additionally, we compare our performance with datafree, which demonstrates that an adversary can successfully steal a model without access to any training data, though this requires a substantial number of queries: 100 queries for graphs of size 250, totaling 25,000 query nodes. We show that our method performs much better in a scenario with very restricted model access (we evaluate our approach on 10, 25, 50, 100, and 500 nodes). We present our results in Table 3 and Appendix Table 5, where we compare all methods using the same surrogate model architecture and different target architectures: SAGE and GAT, with a query limit of 100. We also compare the performance of all methods across different query limits in Figures 4 and 5 (Appendix). Furthermore, in Table 9 (Appendix), we compare the performance using the architectures originally employed in shen2021model and Podhajski_2024. Our method, which does not rely on query embeddings from the victim model, consistently outperforms the previous methods in both accuracy and fidelity. This superiority is evident across various baseline surrogate models, including GIN, GAT, and SAGE. Additionally, by utilizing the randomly initialized encoder, we significantly improved the execution times as seen in Table 9 (Appendix). Moreover, even when the same architecture is used, our approach yields significantly better results, demonstrating its effectiveness.
@@
Transductive setting. For the transductive setting, there is no previous work in the literature to which we can directly compare. However, we believe that in the work by wu2021model we can find a setting to which a comparison of our E2E is reasonable. In particular, unlike our work, wu2021model focuses on the scenarios where the adversary lacks some types of data, e.g., the graph structure or node attributes. In contrast, our focus is on scenarios where the adversary is not limited in terms of data type but in terms of the number of queries that can be made. To create a relevant comparison, we modified the approach by wu2021model by training their model using randomly selected queries under the assumption that it has access to the entire data available to the adversary (with no restrictions on data types). Similarly, we compare our results to datafree, which generates graph queries in a restricted scenario, where the query limit is reduced to just 5, 10, 20, or 50 nodes. The results presented in Tables 4 and 6 (in the Appendix) show a significant improvement in both accuracy and fidelity when our method is applied. Specifically, when the surrogate encoder is trained with all available data and the query selection is optimized using the K-means clustering algorithm, the performance of the stolen model is greatly enhanced. This demonstrates the advantage of our approach in transductive settings, where strategic data utilization and query selection can significantly boost the attack’s effectiveness. We further validate the effectiveness of our method across different surrogate architectures (GIN, SAGE, and GCN) and observe that our method consistently outperforms all baselines regardless of the surrogate used. Notably, the GCN surrogate yields the highest performance overall, surpassing all other methods and architectural configurations. Additionally, Figure 6 shows how the performance improves with different query limits, further emphasizing the impact of our method as the query limit decreases.
@@ 5 Defense
We evaluate all stealing methods under a defense that flips predictions with a probability of in both inductive and transductive settings (Tables 12 and 13 in the Appendix). Our proposed method consistently achieves the highest performance, even in the presence of defense measures. These results highlight the difficulty of defending against our extraction attack, even under strong defense mechanisms that come at the cost of the utility of the model.
@@ B.2 Full Evaluation Results
| Method | Photo | Physics | ||||
| Acc. | Fid. | Rel. Time | Acc. | Fid. | Rel. Time | |
| Target accuracy | 93.0 | 96.0 | ||||
| Method | Physics | CS | Cora | |||
| Acc. | Fid. | Acc. | Fid. | Acc. | Fid. | |
| Farthest-First | 77.04.3 | 77.24.6 | 54.50.2 | 55.00.2 | 60.57.4 | 62.37.1 |
| K-Center Greedy | 77.34.4 | 77.84.5 | 54.50.2 | 55.00.2 | 60.47.4 | 62.47.0 |
| Entropy Sampling | 56.42.3 | 54.12.0 | 57.90.1 | 58.10.1 | 31.87.1 | 34.67.9 |
| Coreset Herding | 88.40.7 | 87.80.5 | 76.10.1 | 76.80.0 | 63.45.4 | 65.44.1 |
| Margin Sampling | 82.70.9 | 83.40.8 | 75.40.3 | 76.10.7 | 56.58.6 | 54.09.7 |
| K-Means (ours) | 91.20.4 | 92.70.5 | 78.42.1 | 79.22.2 | 69.91.2 | 72.51.3 |
| Change | Victim Acc. | Acc. | Fid. |
| Original | 0.96 | 90.00.9 | 84.10.8 |
| Node feature change | |||
| 5% | 0.96 | 90.00.8 | 84.10.8 |
| 10% | 0.95 | 89.60.9 | 83.90.9 |
| 30% | 0.95 | 89.50.6 | 83.80.9 |
| 50% | 0.94 | 88.50.7 | 83.90.7 |
| 70% | 0.92 | 82.82.0 | 80.85.0 |
| Edge dropout | |||
| 30% | 0.95 | 88.50.5 | 89.60.5 |
| 50% | 0.95 | 88.21.0 | 89.11.1 |
| 90% | 0.94 | 81.51.3 | 81.73.0 |
| 100% | 0.95 | 77.32.1 | 77.32.1 |